OpenStudy (xmoses1):

One pipe can fill a tank in 24 minutes, a second can fill it in 8 minutes, and a third can fill it in 12 minutes. If the tank is empty, how long will the three pipes, operating together, take to fill it? 1/6 minute 4 minutes 4 4/5 minutes

OpenStudy (jjuden):

4 min

OpenStudy (jjuden):

Equation: 1/24 + 1/8 + 1/12 = 1/x 1/24 + 3/24 + 2/24 = 1/x 6/24 = 1/x x(1/4) = 1 x = 4

OpenStudy (xmoses1):

Thanks @jjuden

OpenStudy (jjuden):

any more

OpenStudy (anonymous):

^^^ what she said :D

OpenStudy (xmoses1):

No that is it so far :)

OpenStudy (jjuden):

lol @sourwing

OpenStudy (nincompoop):

so do we know what is the capacity of THE tank?

OpenStudy (nincompoop):

a tank and THE tank mean different things a tank is just any thank, THE tank refers to a very specific thank

OpenStudy (anonymous):

we don't need to

OpenStudy (xmoses1):

Please dont lol XP